# Strong constraints on a sharp change in $G$ as a solution to the Hubble tension

### v2

@@ 4 Summary and model extensions
A model along these lines was developed in Desmond et al. (2019), where it was proposed that the Leavitt law in the anchor galaxies NGC 4258, the Milky Way, and the Large Magellanic Cloud is different to that in SN host galaxies due to screening effectively correlating $G$ with one of various measures of environmental density, including local dark matter density (Sakstein et al., 2019). Such a model does not modify SNe themselves, but there is an effective difference between the period–luminosity relations of Cepheids in galaxies with geometric anchors versus SN host galaxies, which are used to calibrate the SN absolute magnitude. This alters the distances to all Hubble flow SNe by the same factor, with corresponding impact on the local $cz^{\prime}$ given the lack of any impact on $z$. Imposing the constraint on $G$ from Cepheid–TRGB consistency, this allows the SH0ES Hubble tension to be reduced to $\approx 2.5\sigma$ (Desmond et al., 2019), while that for the TRGB-calibrated distance ladder is completely resolved (Desmond & Sakstein, 2020). Although it is unclear if this remains the case with the latest measurements from the James Webb Space Telescope (JWST; Freedman et al., 2024; Riess et al., 2024), the screening model appears more attractive than the GSM due to its natural circumvention of the constraints we describe here as well as its far stronger theoretical motivation. Variants of this model are studied in Högâs & Mörtsell (2023a, b).
@@ Acknowledgements
IB and HD are supported by Royal Society University Research Fellowship grant 211046. NS is supported by the AKTION grant number MPC-2023-07755, the Charles University Grants Agency (GAUK) grant number 94224, and the Deutscher Akademischer Austauschdienst Bonn-Prague exchange program. The authors thank Earl Bellinger, Charles Dalang, Hugo Moreira and Jeremy Sakstein for helpful discussions. They also thank the referee for comments which greatly improved this contribution. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@
Appendix A Additional consequences of the GSM & future tests
In this section, we collect a few more qualitative or speculative consequences of the GSM which could be used to construct stringent tests with future data and analysis.
@@ Appendix A Additional consequences of the GSM & future tests
The GSM would also affect the propagation of gravitational waves (GWs; LIGO Collaboration, 2016). It has been argued that GWs would propagate continuously, even during the period in which $G$ varies rapidly (Paraskevas & Perivolaropoulos, 2023). Nonetheless, it seems inevitable that the GSM would have different effects on distances based on stellar populations and those obtained using GWs as standard sirens. Further progress is possible using GWs with an electromagnetic counterpart, with one such instance having been reported so far (GW170817; see Virgo & LIGO Collaborations, 2017). Its distance of 40 Mpc is right where the transition is proposed to have occurred. In the future, it should be possible to find similar events either side of the transition. This might eventually allow a determination of the maximum mass of neutron stars, which is sensitive to $G$ and would therefore differ either side of the transition (Reyes & Sakstein, 2024). The GSM would also lead to a significant change in the gravitational binding energy of any compact object, with neutron stars particularly severely affected (Goldman, 2024).
While the GSM would undoubtedly have severe impacts on the Sun (Section 2.1), it can be easier to model slightly less massive stars, yielding less model-dependent constraints on possible evolution of $G$. Bellinger & Christensen-Dalsgaard (2019) use asteroseismology of the star KIC 7970740 to constrain a model in which $G$ gradually changes as a power-law in time since the Big Bang. The impact would be more severe than the gradual changes over a Hubble time considered by those authors if $G$ were 5% larger than today for nearly the entire 11 Gyr history of KIC 7970740. It would be interesting to constrain the GSM with asteroseismology of this or other stars. However, the age would have to be inferred alongside the other parameters. The Sun offers an advantage in this respect because its age is known from radioactive dating of rock samples.
@@
If the proposed reduction in $G$ occurred on a short enough timescale, any bound orbits would be somewhat destabilised. A sudden drop in the equilibrium circular velocity would cause an object initially on a circular orbit to find itself on a wider elliptical orbit. In the case of the Earth, this situation would arise if the transition occurred over $\ll 1$ yr. Its eccentricity would be ‘pumped’ by a sharp change in $G$. This is disfavoured by the fact that Earth’s orbital eccentricity is only 0.017 and that of Venus is only 0.007.
There would also be consequences elsewhere in the Solar System, including on the giant planets. In general, the whole Solar System would be somewhat destabilised. This would increase the likelihood of giant impacts on the Earth. Indeed, this very possibility has been related to the extinction of the dinosaurs and many other species in the Cretaceous–Paleogene extinction (Perivolaropoulos, 2022), which occurred 66 Myr ago (Renne et al., 2013). This corresponds to a distance of only 20 Mpc, but there are Type Ia SNe with Cepheid calibration out to $\approx 40$ Mpc (e.g., Riess et al., 2022a). Even so, recent studies suggest that the GSM can be reconciled with a transition at a lookback time corresponding to only 20 Mpc (Ruchika et al., 2024a, b). It therefore remains possible that the GSM is associated with the Cretaceous–Paleogene extinction event, though this is disfavoured by geochronometric constraints on the number of days per year a few Myr prior to this very clear geological boundary (see figure 7 of de Winter et al., 2020). It is not presently known whether the asteroid impact largely responsible for it was an isolated incident or part of a more general increase in giant impacts at that time, as would be expected in the GSM.

### v4

@@ 4 Summary and model extensions
A model along these lines was developed in Desmond et al. (2019), where it was proposed that the Leavitt law in the anchor galaxies NGC 4258, the Milky Way, and the Large Magellanic Cloud is different to that in SN host galaxies due to screening effectively correlating $G$ with one of various measures of environmental density, including local dark matter density (Sakstein et al., 2019). Such a model does not modify SNe themselves, but there is an effective difference between the period–luminosity relations of Cepheids in galaxies with geometric anchors versus SN host galaxies, which are used to calibrate the SN absolute magnitude. This alters the distances to all Hubble flow SNe by the same factor, with corresponding impact on the local $cz^{\prime}$ given the lack of any impact on $z$. Imposing the constraint on $G$ from Cepheid–TRGB consistency, this allows the SH0ES Hubble tension to be reduced to $\approx 2.5\sigma$ (Desmond et al., 2019), while that for the TRGB-calibrated distance ladder is completely resolved (Desmond & Sakstein, 2020). Although it is unclear if this remains the case with the latest measurements from the JWST (Freedman et al., 2024; Riess et al., 2024; Jensen et al., 2025), the screening model appears more attractive than the GSM due to its natural circumvention of the constraints we describe here as well as its far stronger theoretical motivation. Variants of this model are studied in Högâs & Mörtsell (2023a, b).
@@ Acknowledgements
IB and HD are supported by Royal Society University Research Fellowship grant 211046. NS is supported by the AKTION grant number MPC-2023-07755, the Charles University Grants Agency (GAUK) grant number 94224, and the Deutscher Akademischer Austauschdienst Bonn-Prague exchange program. The authors thank Earl Bellinger, Charles Dalang, Hugo Moreira, and Jeremy Sakstein for helpful discussions. They also thank the referee for comments which improved this contribution.
@@
Appendix A Additional consequences of the GSM and future tests
In this section, we collect a few more qualitative or speculative consequences of the GSM that could be used to construct stringent tests with future data and analysis.
@@ Appendix A Additional consequences of the GSM and future tests
The GSM would also affect the propagation of gravitational waves (GWs; LIGO Collaboration, 2016). It has been argued that GWs would propagate continuously, even during the period in which $G$ varies rapidly (Paraskevas & Perivolaropoulos, 2023). Nonetheless, it seems inevitable that the GSM would have different effects on distances based on stellar populations and those obtained using GWs as standard sirens. Further progress is possible using GWs with an electromagnetic counterpart, with one such instance having been reported so far (GW170817; see Virgo & LIGO Collaborations, 2017). Its distance of 40 Mpc is right where the transition is proposed to have occurred. In the future, it should be possible to find similar events either side of the transition. This might eventually allow a determination of the maximum mass of neutron stars, which is sensitive to $G$ and would therefore differ either side of the transition (Reyes & Sakstein, 2024). The GSM would also lead to a significant change in the gravitational binding energy of any compact object, with neutron stars particularly strongly affected (Goldman, 2024). Moreover, the entropy of a black hole $\propto G$, so a sharp drop in $G$ seems incompatible with the second law of thermodynamics (van Putten, 2024).
While the GSM would undoubtedly have considerable impact on the Sun (Section 2.1), it can be easier to model slightly less massive stars, yielding less model-dependent constraints on possible evolution of $G$. Bellinger & Christensen-Dalsgaard (2019) use asteroseismology of the star KIC 7970740 to constrain a model in which $G$ gradually changes as a power-law in time since the Big Bang. The impact would be more severe than the gradual changes over a Hubble time considered by those authors if $G$ were 5% larger than today for nearly the entire 11 Gyr history of KIC 7970740. It would be interesting to constrain the GSM with asteroseismology of this or other stars. However, the age would have to be inferred alongside the other parameters. The Sun offers an advantage in this respect because its age is known from radioactive dating of rock samples.
@@
If the proposed reduction in $G$ occurred on a short enough timescale, any bound orbits would be somewhat destabilized. A sudden drop in the equilibrium circular velocity would cause an object initially on a circular orbit to find itself on a wider elliptical orbit. In the case of the Earth, this situation would arise if the transition occurred over $\ll 1$ yr. Its eccentricity would be ‘pumped’ by a sharp change in $G$. This is disfavoured by the fact that Earth’s orbital eccentricity is only 0.017 and that of Venus is only 0.007.
There would also be consequences elsewhere in the Solar System, including on the giant planets. In general, the whole Solar System would be somewhat destabilized. This would increase the likelihood of giant impacts on the Earth. Indeed, this very possibility has been related to the extinction of the dinosaurs and many other species in the Cretaceous–Paleogene extinction (Perivolaropoulos, 2022), which occurred 66 Myr ago (Renne et al., 2013). This corresponds to a distance of only 20 Mpc, but there are Type Ia SNe with Cepheid calibration out to $\approx 40$ Mpc (e.g., Riess et al., 2022a). Even so, recent studies suggest that the GSM can be reconciled with a transition at a lookback time corresponding to only 20 Mpc (Ruchika et al., 2024a, b). It therefore remains possible that the GSM is associated with the Cretaceous–Paleogene extinction event, though this is disfavoured by geochronometric constraints on the number of days per year a few Myr prior to this very clear geological boundary (see figure 7 of de Winter et al., 2020). It is not presently known whether the asteroid impact largely responsible for it was an isolated incident or part of a more general increase in giant impacts at that time, as would be expected in the GSM.